Let the probabilities of binary symbols 0 and 1 be 0.75, 0.25, respectively. Assume no scaling.

a) Determine the arithmetic encoder output bistream for the symbol sequence 0100001.(Determine the tag and truncate it properly)
Answer #1

P(0) = 0.75

P(1) = 0.25

P(0100001) = 0.75 * 0.25 * 0.75 * 0.75 * 0.75 * 0.75 * 0.25 = 0.0148 (ans)